OpenStudy (anonymous):

Choose the two equations you would use to solve the absolute value equation below. Then solve the two equations. |x – 4| = 12

OpenStudy (uri):

|x – 4| = 12 x=12+4 x=16 |x – 4| = 12 x=-12+4 x=-8

OpenStudy (anonymous):

das correct